OpenStudy (anime_lover_levi):

Can some one help me? Medal! Simplify 4sqrt(6)/sqrt(30) by rationalizing the denominator. I know the answer is 4/sqrt(5) but, how do I get to the answer, like what is the math work to get to that answer?

OpenStudy (alones.):

So, basically, it looks like this: http://prntscr.com/b56pdz

OpenStudy (mathstudent55):

How can the answer have a root in the denominator? Aren't you trying to rationalize the denominator?

OpenStudy (anime_lover_levi):

thank you os much alone your a life savor!

OpenStudy (alones.):

No Problem @Anime_Lover_Levi !!

OpenStudy (anime_lover_levi):

I have one more , can y'all help me?

OpenStudy (alones.):

Sure!

OpenStudy (alones.):

Post it in a new post though.

OpenStudy (anime_lover_levi):

ok! (2sqrt(5)+3sqrt(7)^2

OpenStudy (anime_lover_levi):

21+2 sqrt(5) i know thats the answer

OpenStudy (mathstudent55):

\(\dfrac{4 \sqrt 6}{\sqrt {30 }} = \dfrac{4 \sqrt 6}{\sqrt 5\sqrt 6} = \dfrac{4}{\sqrt 5} =\) \(= \dfrac{4 \sqrt 5}{\sqrt 5 \sqrt 5} = \dfrac{4\sqrt 5}{5} \)